# ActiView: Evaluating Active Perception Ability for Multimodal Large Language Models

### v1

@@ 4 Evaluation
To conduct thorough investigation of active perception abilities of MLLMs, we design three evaluation pipelines for different operations of perceptual fields as illustrated in Figure 4, including two fundamental pipelines that apply zooming and shifting respectively, and a mixed pipeline comprising both zooming and shifting operations. We set up five different initial views for each question-image pairs, where a full initial view is used for the zooming and the mixed pipeline, and four constrained views are applied for the shifting pipeline.
Figure 4: Evaluation pipelines for active perception abilities as described in §4.2.
4.1 Models
We investigate both proprietary and open-source models. The proprietary models include widely discussed GPT-4o (OpenAI, 2024), Gemini-1.5-pro (Reid et al., 2024), and Claude 3.5 Sonnet (Anthropic, 2024). For open-source models, we carefully select recent and commonly used models of different structures and of difference scales, such as model families of MiniCPM-V Yao et al. (2024) [^2] https://github.com/OpenBMB/MiniCPM-V, LLaVA (Liu et al., 2023b; c), mPLUG-Owl (Ye et al., 2024b; a), Idefics (Laurençon et al., 2024; 2024), and etc. Since the awareness of fine-grained details and instruction-aware visual features are significant indicators during evaluation, we also include models specifically optimised on these aspects, such as SEAL (Wu & Xie, 2023) for fine-grained details understanding, and Brote (Wang et al., 2024b) which is trained from InstructBLIP (Dai et al., 2023) for instruction-aware and multi-image comprehension. Please refer to Appendix C for details of all 27 evaluated models. These models are divided into two types, single-image models that accepting only one image per input, such as LLaVA-1.6 (Liu et al., 2023b) and MiniCPM-Llama3-V-2.5 (Yao et al., 2024); and multi-image models that allow more than one images to appear in the same input, such as Brote and Idefics. We will describe the approaches for integrating multiple views into the input for the two types of models in § 4.3 and Appendix F.2.
4.2 Evaluation Pipelines
We design three pipelines for different abilities, two fundamental independent abilities, zooming and shifting, and a mixture of them. Generally, the images are split into 4 sub-views to enable zooming and shifting operations, and to ensure that all models are fairly evaluated without bias from their training data or image processing strategies. We also discuss different splittings in Appendix F.1.
Zooming pipeline.
It evaluates the ability to locate and determine fine-grained information that are necessary to answer the question. As illustrated in Figure 4 (a), this pipeline contains two stages, the view selection and the question answering stages. To simulate the zooming operation, models are required to first select sub-views to be zoomed given the initial view, then answer questions based on these zoomed views. The initial view used in this pipeline is the full image with size $w\times h$. Each of the selected sub-views will be resized to size $w\times h$, the same size as the initial view. As demonstrated in Figure 4 (c), the zoomed right-upper view is resized as a $w\times h$ image, and so does the zoomed left-lower view. Afterwards, models answer the question given the two zoomed views. The detailed prompt template are provided in Appendix I.2.
Shifting pipeline.
It evaluates the ability to shift perceptual fields for missing information and to deduce the answer given perceived perceptual fields following templates in Appendix I.3. This is also a two-stage pipeline as shown in Figure 4 (b). To simulate the movement of human eyes, models are presented with an initial view, of size $w\times h$, which is a cropped field from the original image, and are asked to determine if the current view or views are sufficient for answering the questions. Upon receiving positive responses, models are prompted to produce answer given the current view or views. If the model requires more views to infer the answer, an adjacent view will be given until the model is able to answer the question. For this pipeline, we assign different difficulties according to human-annotated visual clues, described in § 3.2, contained in the initial views as follows:
  • Shifting-R: starting the missing view examination with randomly selected initial views.
  • Shifting-E: easy-level evaluation, where initial views contain at least one entire visual clue for answering the question.
  • Shifting-M: medium-level evaluation, where initial views contain only partial visual clues for answering the question.
  • Shifting-H: hard-level evaluation, where no visual clues appear in the initial views.
Mixed pipeline.
While the above pipelines allow for either zooming or shifting individually, we also implement a mixed setting that does not specify the type of active perception ability required. As illustrated in Figure 4 (c), models must independently decide whether to zoom and/or shift to different perceptual fields. Unlike the zooming pipeline, where the model answers questions based on all selected views, in the mixed pipeline, a view would be discarded after zooming if the model recognizes it as irrelevant to the question. In comparison to the shifting pipeline, the mixed pipeline also provides access to the full image view in addition to cropped sub-views. Please refer to Appendix I.4 for employed prompt templates. This type of evaluation requires models to account for all the sub-views and the full image for unbiased operation determination and view selection. Otherwise, it is at risk of reverting to either zooming or shifting evaluation without sufficient and unconverted visual information. Therefore, the mixed pipeline is only applied to multi-image models during evaluation.
In addition to these pipelines for evaluating active perception abilities, our benchmark also supports the general VQA evaluation, where models are prompted to answer visual questions given original images without zooming or shifting. The detailed prompt templates are provided in Appendix I.1. We also evaluate the performance of this setting to assess the difficulty of our created benchmark.
Table 2: Results of individual abilities required by active perception, following shifting and zooming pipelines described in §4.2. We report the results of 20 widely-discussed models in this table, and refer readers to Table 7 for results of 27 evaluated models. The humane performance is 84.67% referring to Table 5 in Appendix B. “Model AVG”: average scores of zooming (column “Zooming”) and shifting (columns “Shifting-R”, “Shifting-E”, “Shifting-M”, and “Shifting-H”) evaluations. The best scores of each column are bolded and the best scores in each model types are highlighted.
Models Zooming Ability Shifting Ability Models AVG
Full image Zooming Single View Shifting-R Shifting-E Shifting-M Shifting-H
Proprietary models (APIs)
Gemini-1.5-pro 73.85 72.31 58.15 67.08 67.38 65.54 67.69 68.00
GPT-4o 67.38 68.62 61.23 67.08 66.77 65.23 64.31 65.54
Claude 3.5 Sonnet 72.92 71.69 54.46 65.23 66.15 60.31 61.85 65.05
Open-source models for multiple images as input
Qwen2-VL 63.08 64.62 54.46 61.23 62.77 64.31 61.85 62.96
Idefics3-8B-Llama3 59.08 58.15 53.23 61.85 59.38 59.69 60.31 59.88
MiniCPM-V 2.6 64.62 61.85 54.46 54.77 61.23 58.15 55.69 58.34
mPLUG-Owl3 62.46 60.92 54.15 51.69 56.31 55.69 53.54 55.63
LLaVA-OneVision 64.92 65.23 56.92 53.54 57.23 52.31 48.62 55.39
InternVL2-8B 58.15 56.00 45.85 54.77 59.70 53.23 52.00 55.14
idefics2-8b 61.85 61.85 55.69 53.23 56.92 51.69 49.23 54.58
Mantis 59.08 60.62 52.92 52.92 55.38 52.92 52.31 54.83
Brote-IM-XL-3B 54.77 54.46 55.69 51.38 51.08 52.62 47.69 51.45
MMICL-XXL-11B 51.69 49.54 50.15 49.85 49.85 46.77 45.54 48.31
@@ 4.2 Evaluation Pipelines
Open-source models for single image as input
GLM-4V-9B 67.08 56.92 53.85 56.92 60.62 56.00 52.92 56.68
InternVL-Vicuna-13B 56.92 62.77 52.31 53.85 52.92 52.92 51.08 54.71
LLaVA-1.6 7B 55.08 68.92 50.15 51.69 52.31 49.23 48.00 54.03
mPLUG-Owl2-7B 55.08 55.38 52.00 47.38 46.46 46.46 46.15 48.37
Mini-Gemini-7B-HD 55.69 34.77 51.70 48.62 48.00 47.69 50.15 45.85
SEAL 48.31 54.77 42.77 42.15 42.77 40.02 40.62 44.07
Table 3: Results of mixed pipeline for multi-image models. “ACC”: accuracy. “#zoom”: average zooming operations. “#shift”: average shifting operations. “#view”: average used views.
GPT-4o Qwen2-VL MiniCPM-V 2.6 mPLUG-Owl3 Idefics3
ACC #zoom #shift #view ACC #zoom #shift #view ACC #zoom #shift #view ACC #zoom #shift #view ACC #zoom #shift #view
69.54 1.61 1.23 1.35 65.54 2.51 2.17 2.12 64.00 1.31 0.39 0.94 59.69 2.59 1.49 1.43 62.15 1.16 0.59 0.58
4.3 Processing of Views
The question answering stage of the zooming, shifting, and mixed pipelines, as well as the missing view examination stage of the shifting pipeline, require multi-image inputs if multiple views are selected. In this paper, we primarily focus on the interleaved multi-image setting, since it is more practical and natural compared to the single-image setting. Multi-image models can naturally read and understand multiple views at one time (in the form of different images) during evaluating, and we directly format the images and text in an interleaved format. However, we also propose methods for evaluating powerful single-image models. For these models, we employ two strategies to enable simultaneous understanding of different views. One is to concatenate the required views into a single flattened image, and the other preserves merely the current view as an image while converting the remainings into textual descriptions. Please refer to Appendix F.2 for details.
5 Results and Analysis
Main results. We adopt accuracy as the evaluation metric for question answering. Experimental results of zooming and shifting pipelines on our benchmark are listed in Table 2, whose elaborated results on each type are provided in Appendix E. Results of human evaluations are discussed in detail in Appendix B. We obtain the random choice result of 33.95%, averaged over 10k runs.
Firstly, we can conclude from Table 2 that all the evaluated models perform better than random guessing, indicating their potential to maintain active perception abilities of zooming and shifting. However, even the average results of powerful proprietary models are still much lower then human performance (84.67%, as in Table 5). Secondly, although proprietary models achieve better overall performances compared to open-source models, the performance gap between these two categories are considerably smaller compared to gaps observed in other tasks from previous research. With our evaluation pipelines, proprietary models achieve the highest average score of 68.00%, whereas the highest from open-source models is 62.96%. Thirdly, among open-source models, multi-image models present better average results compared to single-image models, particularly in shifting evaluations where models can only access constrained views.
For results of mixed evaluation in Table 3, we notice that the evaluated models benefit from enabling complex active perception and outperform individual zooming or shifting settings on average. Among the evaluated models, MiniCPM-V 2.6 (64.00%) and Idefics3-8B-Llama3 (62.15%) achieve even higher accuracy compared to the results of providing human-annotated ground truth views from Table 4 (62.77% and 60.92%, respectively). The mixed pipeline encourages models to zoom and/or shift perceptual fields autonomously, similar to human behavior, and these results demonstrate the significance and effectiveness of active perception abilities. However, during experiments, we observe that some multi-image models fail to follow instructions of mixed evaluation by generating irrelevant responses or selecting invalid views, which, unfortunately, disrupts the mixed process.
@@ 5 Results and Analysis
Multi-image Models
GPT-4o 73.54 68.62 69.03 2.29 67.08 60.54 3.26
mPLUG-Owl3 60.62 60.92 68.57 2.66 51.69 74.62 4.00
Claude 3.5 Sonnet 72.31 71.69 67.64 2.19 65.23 45.52 2.47
Qwen2-VL 65.85 64.62 64.61 2.35 61.23 74.62 4.00
Gemini-1.5-pro 72.00 72.31 62.63 2.10 67.08 46.33 2.46
MiniCPM-V 2.6 62.77 61.85 57.03 2.20 54.77 54.83 2.98
LLaVA-OneVision 64.92 65.23 46.67 2.35 53.54 37.14 2.02
Idefics3-8B-Llama3 60.92 58.15 41.09 1.52 61.85 74.62 4.00
InternVL2-8B 73.23 56.00 41.09 1.53 54.77 45.75 2.61
Single-image Models
InternVL-Vicuna-13B 68.00 62.77 83.47 3.31 53.85 69.73 3.75
LLaVA-1.6 13B 67.69 68.92 68.57 2.65 51.69 74.62 4.00
SEAL 56.92 54.77 68.22 2.74 42.15 71.48 3.83
MiniCPM-Llama3-V-2.5 62.20 61.25 66.12 2.46 53.85 63.56 3.42
mPLUG-Owl2-7B 67.38 55.38 47.61 1.97 47.38 74.62 4.00
GLM-4V-9B 74.46 56.92 30.62 1.08 56.92 50.17 2.64
Impacts of selected views.
@@
Our evaluation pipelines involve selecting useful view in their first stages. The reliability of selected views plays a crucial role in the following question answering stage. We compute the recall of used views following Equation 1 in Appendix D, and include results in Table 4, along with accuracy of providing models with groundtruth views. Overall, higher selection recall tends to correlate with higher question answering accuracy. For example, Idefics3-8B-Llama3 and InternVL2-8B present the lowest recalls (41.09%) among multi-image models in Table 4, leading to lower zooming evaluation accuracies of 56.00% and 58.15%, respectively. For the shifting evaluation, some models keep shifting view until all four views are inquired. However, this does not necessarily support a better accuracy, as some views contain redundant information that might distract the model during reasoning. Additionally, we observe shifting evaluations tend to require more views to be used for answering questions than zooming evaluations, yet it often results in inferior overall performance compared to zooming. This is because some of the current advanced models struggle to either move their field of views for necessary visual details, or screen out distracting information. Therefore, we believe that more attention should be paid to evaluating and enhancing active perception abilities of MLLMs given constraint perceptual fields.
Analysis of difficulties for shifting evaluation.
@@
Generally, the accuracy of question answering and the recall of view selection decrease as the difficulties of the initial views increases. As shown by typical results for LLaVA-OneVision and GLM-4V-9B in Table 2, the gaps between shifting-E and shifting-H are as large as 8.61% and 7.70%, respectively. However, we observe some exceptions from Gemini-1.5-pro and Idefics3, where the accuracy of shifting-H is slightly higher than shifting-E. For Gemini-1.5-pro, the recall for shifting-H is 47.73%, over 1 point higher than shifting-E (45.29%). We conclude that Gemini-1.5-pro achieves higher accuracy on shifting-H due to the acquisition of more proper views. However, Idefics3 presents a different trend. It maintains a recall of 74.64% from shifting-E to shifting-H, but achieves a higher accuracy on shifting-H. We hypothesis that the performance gain of this model comes from the order of input views, where hard-level evaluation starts with less relevant views and appends more useful views at the end of the image sequence when all the views are selected.
Analysis of view splitting and image processing strategies.

### v2

@@ 4 Evaluation
For thorough investigation, we design three evaluation pipelines for different operations of perceptual fields as illustrated in Figure 4, including two individual pipelines for core components, and a mixed pipeline incorporating both. We set up five different initial views for each question-image pairs, where a full image of limited resolution is used for zooming and mixed pipelines, and four constrained views are applied for the shifting pipeline. These correspond to the 1,625 evaluation instances mentioned in previous section. Due to page limitation, details of pipelines can be found in Appendix E, and evaluated models are discussed in Appendix D.
@@
Metrics Proprietary Models Multi-image Open-source Models Single-image Open-source Models
Gemini
1.5-pro
GPT-4o
Claude3.5
Sonnet
Qwen2.5
VL-7B
DeepSeek
-VL2
Idefics
-3-8B
MiniCPM
-V 2.6
mPLUG-
Owl3-7B
LLaVA
-OV
Intern
VL2-8B
Mantis
-8B
Phi-3.5
-vision
GLM-
4V-9B
InternVL
13B
LLaVA
-1.6 7B
MGM
-7B-HD
SEAL
Zooming $P_{\text{select}}$ 79.00 79.60 81.49 73.72 71.64 71.31 69.13 68.25 69.08 71.03 64.88 69.60 74.93 70.48 65.46 66.24 65.92
$R_{\text{select}}$ 62.63 69.03 67.64 70.55 55.30 41.09 57.03 68.57 46.47 41.09 22.80 28.52 30.62 61.70 68.57 30.15 68.22
$F_{1}$ 69.87 73.94 73.92 72.10 62.42 52.14 62.50 68.41 47.91 52.06 33.74 40.46 43.47 65.80 66.98 41.44 67.05
@@
[…]
Average ACC 68.00 66.40 65.05 68.00 65.11 59.88 58.34 55.63 55.39 55.14 54.83 51.38 56.68 54.71 54.03 45.84 44.07
ACC w/
Human Clues
72.00 73.54 72.31 70.77 68.31 60.92 62.77 60.62 64.92 73.23 59.38 58.15 74.46 68.00 67.69 68.31 56.92
Table 2: Results of evaluation of individual components, following shifting and zooming pipelines. We list results of some widely-discussed models here, and refer readers to Table 11 for more details. The human performance is 84.67% referring to Table 6 in Appendix B. “Average AVG”: average scores of question answering accuracy of all settings. The best scores of each row are bolded and the best scores in the other model types are highlighted.
Metrics Proprietary Models Multi-image Open-source Models
Claude 3.5
Sonnet
GPT-4o Gemini-1.5-pro
Qwen2.5-VL
-7B
Qwen2-VL
-7B
DeepSeek-VL2
Qwen2.5-VL
-3B
MiniCPM
-V 2.6
Idefics3
-8B
mPLUG-Owl3
-7B
#zoom 2.30 1.61 1.82 1.88 2.51 2.65 1.21 1.31 1.16 2.59
#shift 3.15 1.23 1.65 1.59 2.17 1.74 1.73 0.39 0.59 1.49
#view (diff) 1.89 (-0.75) 1.35 (-1.29) 1.14 (-1.50) 1.26 (-1.38) 2.12 (-0.52) 2.43 (-0.21) 0.94 (-1.70) 0.94 (-1.70) 0.58 (-2.06) 1.43 (-1.21)
ACC 72.00 69.54 68.92 70.77 65.54 65.23 64.62 64.00 62.15 59.69
Table 3: Experimental results of mixed pipeline for integrated components. “#zoom”: average zooming operations; “#shift”: average shifting operations. The most count is marked in blue and the least count in yellow. “#view”: average used views; “diff”: #view-#view_annotated, indicating the difference between actual selected views and the average sub-views containing visual clues (2.64). “ACC”: question answering accuracy. The best results for #view (diff) and ACC are bolded and the second best are underlined.
4.1 Pipelines for Individual Component
We separately investigate two core components, zooming and shifting. The zooming pipeline evaluates the ability to locate and determine necessary fine-grained information. As shown in Figure 4 (a), this pipeline contains two stages, view selection and question answering. In this pipeline, models first select sub-views to zoom in given the initial view, the full image of size $w\times h$, then answer the question based on these zoomed views. The selected sub-views are resized to $w\times h$, the same as the initial view, to enable a zooming operation. Note that a “None” selection is permitted.
The shifting pipeline addresses the ability to navigate perceptual fields incrementally, mimicking real-world scenarios where complete context is unavailable. It measures the ability to shift perceptual fields for missing information and to infer the answer jointly based on constrained perceptual fields. This is also a two-stage pipeline as in Figure 4 (b). To simulate the movement of human eyes, the model begins with an initial view of size $w\times h$, and determines if the current views are sufficient for answering the question. If more views are needed, adjacent views will present until the answer can be inferred. Furthermore, we assign four difficulty levels according to human-annotated visual clues contained in the initial views, namely “Shift-R”, “Shift-E”, “Shift-M”, and “Shift-H”, with corresponding settings are detailed in Appendix E.
4.2 Pipeline for Integrated Components
In addition to the individual pipelines, we also implement an automated mixed setting, mixed pipeline, that does not specify the type of active perception ability required. As illustrated in Figure 4 (c), models must independently decide whether to zoom, shift, or use both to address different perceptual fields. In contrast to the zooming pipeline, where models answer questions based on all selected views, the mixed pipeline allows models to discard irrelevant views after selection. Unlike the shifting pipeline, the mixed pipeline also provides access to the full image view in addition to cropped sub-views. The mixed pipeline emphasizes model autonomy and requires models to account for all views, including the full one, to ensure unbiased operation decision and view selection. Otherwise, it is at risk of reverting to either zooming or shifting tasks, limiting its evaluation. The need for autonomy, strict adherence to instructions, and comprehensive understanding of all views makes this pipeline suitable only for the most advanced and recent multi-image models.
Templates for the above pipelines are listed in Appendix J.2,  J.3 and  J.4, respectively. In addition, we enable a general VQA evaluation, where models answer visual questions given full images without zooming or shifting, to serve as the reference for assessing the difficulty of our created benchmark. Detailed prompts are provided in Appendix J.1.
4.3 Processing of Views
In this paper, we focus on the interleaved multi-image setting, which is more practical and natural than the single-image setting. Multi-image models can naturally comprehend several views at one time during evaluating, allowing us to directly format the images and text in interleaved form. For fairness, we also design evaluation methods for single-image models. To enable zooming and shifting operations, images are split into four sub-views, ensuring unbiased evaluation across all models, regardless of their training data or image processing strategies. We also discuss different splitting methods (Appendix F.1), and input view processing techniques, including image processing (Appendix F.2) and textual form conversion (Appendix F.3).
5 Results and Analysis
- Human* Random Text-only (GPT-4o)
ACC 84.67 33.95 2.45
Table 4: Human level performances, random choice result (averaged over 10k runs), and text-only evaluation. Detailed discussion of human evaluation is provided in Appendix B, and text-only evaluation across different models is provided in Appendix C.
Experimental results of individual components, zooming and shifting, are in listed Table 2. We adopt accuracy as the evaluation metric for question answering, together with measurements of view selection (detailed in Appendix G). Due to space constraints, we only list selected high-performance models in these two tables, and provide elaborated results of all evaluated models in Appendix H, including scores for each categories, sub-classes and difficulties. Experimental results of integrated components from the mixed pipeline are reported in Table 3. The number of candidate options ranges from two to seven, with a random choice baseline of 33.95% on our benchmark. We also conduct human and text-only evaluation to assess the difficulty and robustness of our benchmark. The average performance among six testers is 84.67%, indicating that while our benchmark is feasible for humans, it can still be challenging. The text-only evaluation implies that questions in ActiView cannot be solved solely by commonsense knowledge within models, and that visual information is crucial for completing the tasks.
5.1 Main Results
Results of evaluation for individual components.
We draw four key findings from the pipelines for individual components of active perception.
First, as shown in Table 2, all evaluated models outperform random guessing, indicating their potential to maintain active perception abilities of zooming and shifting. However, even the best proprietary models fall significantly behind human. Second, although proprietary models achieve better overall performances compared to open-source models, the performance gap between these two categories are considerably smaller compared to gaps observed in other tasks from previous research. Moreover, the gap is becoming smaller for some recent released open-source models such as Qwen2.5-VL, which achieves the same highest average score of 68.00% as Gemini-1.5-pro. Third, among open-source models, multi-image models largely outperform single-image models, particularly in shifting evaluations with constrained views. Finally, we observe that the view selection scores, F1 in particular, are highly related to the final performance. Lower scores of precision, recall and F1 stand for more unnecessary information given, which also correlates with lower QA performances.
For results of mixed evaluation in Table 3, we observe that the evaluated models benefit from enabling complex active perception and often outperform individual zooming or shifting on average. Notably, MiniCPM-V 2.6 (64.00%) and Idefics3-8B-Llama3 (62.15%) surpass the accuracy of given human-annotated views (62.77% and 60.92%, respectively, from Table 17); and Claude 3.5 Sonnet and Qwen2.5-VL-7B achieve equivalent performance of given human-annotated visual clues.
Results of evaluation for integrated components.
The mixed pipeline encourages models to zoom and/or shift perceptual fields autonomously, mimicking human behaviors and highlighting the effectiveness of active perception. However, during experiments, we noticed that some large multi-image models, typically released months or a year ago, failed to follow instructions in the mixed evaluation, generating irrelevant responses or selecting invalid views, disrupting the mixed process. In contrast, recent models, regardless of size, successfully handled the this evaluation. With the reported counts of operations, we can conclude that models that actively zoom and shift views are likely to present higher question answering scores, thereby exhibiting better active perception ability.
Models W/ Clues Full Zooming
GPT-4o 73.54 67.38 68.62
Qwen2-VL-7B 65.85 63.08 64.62
Qwen2.5-VL-3B 66.15 65.85 66.15
Qwen2.5-VL-7B 70.77 67.08 68.92
InterVL2-8B 73.23 58.15 56.00
Qwen2-VL-7B 65.85 63.08 64.62
Idefics3-8B 60.92 59.08 58.15
SEAL 56.92 48.31 54.77
Table 5: Performance comparison among providing annotation clues (W/ Clues), full images without applying perceptual constraints (Full), and our designed zooming pipeline (Zooming).
Discussion of the necessity of active perception.
To address the effectiveness and significance of active perception, we conduct comparison among three settings: i) providing annotation clues, ii) full images without applying perceptual constraints, and iii) our designed zooming pipeline, to demonstrate the usefulness of active perception, and also highlight the current limitations of models in active perception. We select several high-performance models from our experiments, and report corresponding results in Table 5, where active perception improves over Full for both manually guided scenarios (W/ Clues) and model automated scenarios (Zooming). Results suggest that if models could accurately identify the necessary views (such as obtaining the human-annotated clues), their performance could be further improved via active perception. More case studies are further discussed in Appendix I to support the necessity of active perception.
5.2 Analysis of Different Pipelines
Impact of selected views
Our pipelines involve selecting useful view in their first stages. The reliability of selected views plays a crucial role in the following question answering stage. We refer readers to Appendix H.3 for elaborate discussion on corresponding results (in Table 17).
@@ 5.2 Analysis of Different Pipelines
Overall, lower selection recall tends to correlate with lower VQA accuracy. For example, Idefics3-8B-Llama3 and InternVL2-8B present the lowest recalls (41.09%) among multi-image models, leading to lower accuracies for zooming evaluation, 56.00% and 58.15%, respectively. We also investigate the performance when given groundtruth views that contain human-annotated clues. Generally, models are prompted to generate more accurate answers compared to the pure zooming setting. However, mPLUG-Owl3, Gemini-1.5-pro, and LLaVA-OneVision are only exceptional, whose performance slightly decrease when given visual clues. We argue that they are better at the question answering task rather than exhibiting active perception ability. Additionally, we observe that shifting evaluations tend to require more views for answering questions compared to zooming evaluation, yet often results in inferior overall performance, indicating that models lack the ability to actively shifting perceptual fields under constraints. Thus, we believe that more attention should be paid to evaluating and enhancing active perception abilities of MLLMs given limited perceptual fields.
Performance for different difficulty levels
@@
Generally, the accuracy of question answering and the recall of view selection decrease as the difficulties of the initial views increases. As shown by typical results of LLaVA-OneVision and GLM-4V-9B in Table 2, the gaps between easy and hard settings are as large as 8.61% and 7.70%, respectively. However, exceptions exist for Gemini-1.5-pro and Idefics3, demonstrating different reasons, where one is caused by the recall of selected views, and the other lies in the order of relevant views. Gemini-1.5-pro presents higher recall on Shift-H due to higher selection recall. Idefics3 maintains the same recall for all different settings, but achieves a higher accuracy on Shift-H. We hypothesis that the gain comes from the order of input views, where hard-level evaluation starts with less relevant views while appends more informative views at the end of input image sequence when all the views are selected. Please refer to Appendix H.2 for detailed analysis on shifting evaluation.
5.3 Analysis of View Processing Strategies